if the body mass is 5 kgs .the volume is 4.0 cm3.then what is the density?

Grade:11

2 Answers

pavan kumar
10 years ago
1.25kg/cm3
Harsh Srivastava
10 years ago
1 cm3=10^-6 m3 Hence 4 cm3 = 4*10^-6 m3 Density = Mass/Volume = 5/4*10^-6 =1.25*10^6 kg/m3